Need help with both questions!

1.) According to data recorded over many years, 10% of the population of adult females is left-handed. A tennis player wants to test if the proportion is different for professional women's tennis players. Out of a random sample of 100 professional women's tennis players, 14 were left-handed.

The sample data can be used to perform a two-sided hypothesis test to test the claim that the proportion of left-handed professional women's tennis players is not equal to the proportion for all adult females.

Based on the given information and using the appropriate formula, calculate the test statistic for this hypothesis test. Round your answer to three decimal places. Enter the numeric value of the test statistic in the space below.

2.) According to data recorded over many years, 10% of the population of adult females is left-handed. A tennis player wants to test if the proportion is different for professional women's tennis players. Out of a random sample of 100 professional women's tennis players, 12 were left-handed.

The sample data can be used to perform a two-sided hypothesis test to test the claim that the proportion of left-handed professional women's tennis players is not equal to the proportion for all adult females.

Based on the given information and using the appropriate formula, calculate the test statistic for this hypothesis test. Round your answer to three decimal places. Enter the numeric value of the test statistic in the space below.
Answer #1

1)

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: p = 0.1
Alternative Hypothesis, Ha: p ≠ 0.1

Test statistic,
z = (pcap - p)/sqrt(p*(1-p)/n)
z = (0.14 - 0.1)/sqrt(0.1*(1-0.1)/100)
z = 1.333

2)


Null Hypothesis, H0: p = 0.1
Alternative Hypothesis, Ha: p ≠ 0.1

Test statistic,
z = (pcap - p)/sqrt(p*(1-p)/n)
z = (0.12 - 0.1)/sqrt(0.1*(1-0.1)/100)
z = 0.667